At an output level of 49,000 units, you calculate that the degree of operating leverage is 3.60. Suppose fixed costs are $250,000. What is the operating cash flow at 43,000 units? (Do not round intermediate calculations and round your final answer to the nearest whole number, e.g., 32.) Operating cash flow $ What is the new degree of operating leverage? (Do not round intermediate calculations and round your final answer to 3 decimal places, e.g., 32.161.) Degree of operating leverage
Answer #1

A)Contribution margin at 43000 units = 43000 *18.36735

                            = 789796.05

Operating cash flow =contribution margin -fixed cost

                        = 789796.05 -250000

                        = 539796.05   (rounded to 539,796)

Working:

Degree of Operating leverage(at 49000 units) =contribution margin /Fixed cost

          3.60 = Contribution margin / 250000

        Contribution margin = 250000 *3.60

                            = 900000

contribution per unit =Total contribution /number of units

                         = 900000/49000

                         = 18.36735

Fixed cost remains constant .

b)New degree of operating leverage= contribution margin /net operating cash flow

                                         = 789796.05 / 539796.05

                                        = 1.463